OpenStudy (anonymous):

I Need help figuring this one out...not sure where to start with it. Find the equation of a horizontal line that contains (-6,8)

hartnn (hartnn):

Equation of horizontal line is in the form of y=c what is y-co-ordinate in (-6,8) ?

hartnn (hartnn):

your name is awesome! and funny :P

OpenStudy (anonymous):

8 right?

OpenStudy (anonymous):

haha thanks

hartnn (hartnn):

right! so the equation of horizontal line will be y=8 because (-6,8) has to lie on it.

OpenStudy (anonymous):

this is what it says under the question... does it still make it y=8?

OpenStudy (anonymous):

Think about what the slope is of a horizontal line and use that slope and the point and place into point-slope and then transform to slope-intercept form.

hartnn (hartnn):

OMG! ok, what do u think is the slope of horizontal line is ?

OpenStudy (anonymous):

zero slope

OpenStudy (anonymous):

so i do 0x +8y=

hartnn (hartnn):

it said ' place into point-slope' so, put m=0 in y-y1= m(x-x1)

OpenStudy (anonymous):

oh i am still figuring out the slope formulas. my final exam is today and im still really lost on this stuff! :(

hartnn (hartnn):

slope = m = 0 is correct. just put it in y-y1= m(x-x1) what u get ?

OpenStudy (anonymous):

i got y=-6

hartnn (hartnn):

y-y1 = 0(x-x1) y-y1=0 ok?

hartnn (hartnn):

now put point (x1,y1) = (-6,8) so put y1=8 in that

OpenStudy (anonymous):

right... i did y--6=0(x--8) which comes out to be y+6=0 so subtract 6 from both sides and y=-6 right? or no?

hartnn (hartnn):

x1=6, y1=8 not the other way round

OpenStudy (anonymous):

oooops!!!!! okay ill do it again

OpenStudy (anonymous):

so y=8

hartnn (hartnn):

yes

OpenStudy (anonymous):

haha so why did i have to do all that if i know what each coordinate is? shes just making us do extra work or what?

hartnn (hartnn):

she's giving u a particular way to solve, what i did in 1st comment was a shortcut., never do that.

OpenStudy (anonymous):

oh okay. you are so awesome, thank you!! :)

hartnn (hartnn):

lol, welcome ^_^

OpenStudy (anonymous):

how do i fugire this one out? same kinda thing but with x?

OpenStudy (anonymous):

Find the equation of a vertical line that contains (-9,4) Think about what the slope is of a vertical line and if the equation will start with x or y=

hartnn (hartnn):

vertical line slope = ?

OpenStudy (anonymous):

i dont know... -9?

OpenStudy (anonymous):

undefined

hartnn (hartnn):

un defined, so the equation is in the form of x=c

hartnn (hartnn):

whats x-cordinate ?